# On Nontrivial Winning and Losing Parameters of Schmidt Games

### v1

@@ 6. B-Badly Approximable Numbers
    $\left(\frac{1}{1-\beta}\right)\leq B^{N}|A_{t}|\leq\left(\frac{1}{1-\beta}\right)(1+\varepsilon)=\left(\frac{2c(1-\alpha\beta)}{\beta\gamma_{A}}\right),$ (6.2)
    and since $|A_{t}|(1-\beta)>1/B^{N}$, Bob can play his next interval centered at $m/B^{N}$ for some $m\in\mathbb{Z}$. Furthermore, by Lemma 1.2, Bob can play in such a way that in finitely many moves, $B_{t+s}\subset J_{n}$. By the Equidistribution Theorem, there exists an infinite number of possible $t$ that satisfy inequality 6.2. Therefore, Bob can do this an infinite number of times, and Alice loses.
Theorem 6.5.
The set $\operatorname{\widetilde{2-BA}}(c)$ is winning for $\alpha<\frac{1}{3}$, $\beta>\frac{9c}{1-3\alpha}$, and $c<1/6$.
Proof.
The winning strategy for Alice is as follows: suppose Alice is given an interval $B_{i}$ of length $l_{i}$. Pick $k_{i}$ such that
$\frac{1}{2^{k_{i}}}<l_{i}\leq\frac{1}{2^{k_{i}-1}}$
or equivalently,
$\frac{l_{i}}{2}\leq\frac{1}{2^{k_{i}}}<l_{i}.$
Let $W_{i}$ denote the set of all intervals of length $\alpha l_{i}$ and center $c_{i}$ of the form $(m\pm\frac{1}{3})2^{-k_{i}}$, where $m\in{\mathbb{Z}}$. Then, Alice will select an interval from $W_{i}$. This is always possible because the maximum distance between the two endpoints of sets is
$\left|c_{i}-c^{\prime}_{i}\right|+\alpha l_{i}=\frac{1}{3\cdot 2^{k_{i}-1}}+\alpha l_{i}<\frac{2l_{i}}{3}+\frac{l_{i}}{3}=l_{i}$
and thus any interval of length $l_{i}$ must contain at least one of the intervals in $W_{i}$.
Next, define $q_{i}$ to be the smallest integer such that $m/2^{q_{i}}\in A_{i}$ for some $m\in{\mathbb{Z}}$. Note that
$\left|\frac{m}{2^{q_{i}}}-c_{i}\right|=\frac{1}{3\cdot 2^{q_{i}}}<\frac{\alpha l_{i}}{2}<\frac{1}{3\cdot 2^{q_{i}-1}}.$
@@
$\left|x-\frac{m}{2^{n}}\right|>\frac{c}{2^{n}}\ \text{for all}\ n\leq q_{i}-2.$
by induction on $i$. Suppose that the statement holds for $i$. Then, Bob picks an interval $B_{i+1}$ with length $l_{i+1}=\alpha\beta l_{i}$. Now, note that for all $x\in A_{i+1}$ and $q_{i}\leq n\leq k_{i+1}$,
@@
$>\frac{c}{2^{q_{i}}}\geq\frac{c}{2^{n}}.$
To take care of $k_{i+1}+1\leq n\leq q_{i+1}-2$, simply note that for all $x\in A_{i+1}$ and $c<1/6$,
$\left|x-\frac{m}{2^{n}}\right|$ $>\frac{1}{3\cdot 2^{n}}-\frac{\alpha l_{i+1}}{2}>\frac{1}{3\cdot 2^{n}}-\frac{1}{3\cdot 2^{q_{i+1}-1}}>\frac{c}{2^{n}}.$
Finally, for $n=q_{i}-1$, we split into cases based on $A_{i}$, the prior move.
@@
    then for all $x\in A_{i}$,
    $\left|x-\frac{m}{2^{q_{i}-1}}\right|$ $\geq\frac{1}{3\cdot 2^{q_{i}-1}}-\frac{\alpha l_{i}}{2}$
    $>\frac{1}{3\cdot 2^{q_{i}-1}}-\frac{1}{3\cdot 2^{q_{i}-1}}+\frac{c}{2^{q_{i}-1}}$
    $=\frac{c}{2^{q_{i}-1}}.$
@@
    Then, if
    $B_{i+1}\cap\left[\frac{m-c}{2^{q_{i}-1}},\frac{m+c}{2^{q_{i}-1}}\right]=\emptyset,$
    it is clear that Alice does not have to worry about $n=q_{i}-1$, since for all $x\in A_{i+1}\subset B_{i+1}$, $|x-m/2^{q_{i}-1}|>c/2^{q_{i}+1}$. Thus, for the remainder of this case, assume $B_{i+1}$ intersects the interval. Without loss of generality, assume that $m/2^{q_{i}-1}>x$ for all $x\in B_{i+1}$ (that is, the fraction lies to the right of the interval $B_{i+1}$). This is possible because the mapping $x\mapsto-x$ gives us the other case. Alice will then choose an interval $A_{i+1}\in W_{i+1}$ centered at $c_{i+1}$ such that
@@
    that is, Alice cannot choose the closest interval to $\frac{m}{2^{q_{i}-1}}$. This is always possible because the second closest interval is within $l_{i+1}$ of $\frac{m}{2^{q_{i}-1}}$; that is, for $x\in A_{i+1}$,
    $\left|x-\frac{m}{2^{q_{i}-1}}\right|<\frac{2}{3}\left(\frac{1}{2^{k_{i+1}}}\right)+\frac{\alpha l_{i+1}}{2}<\frac{2}{3}l_{i+1}+\frac{1}{6}l_{i+1}<l_{i+1}$
    and thus whenever Alice is able to pick the closest interval, Alice is also able to pick the second closest interval. Furthermore, for $x\in A_{i+1}$,
@@
    $>\frac{c}{2^{q_{i}-1}}$
    proving this case and completing the induction.
Thus, $x_{\infty}\in\operatorname{\widetilde{2-BA}}(c)$ and Alice wins.
A special case of winning and losing zones given by Theorem 6.4 and Theorem 6.5 can be seen on Figure 5.
We also consider the special case of $\widetilde{\text{3-BA}}(1/6)$, defined as
$\widetilde{\text{3-BA}}(1/6):=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{3^{k}}\right|\geq\frac{1}{6\cdot 3^{k}}\text{ for all }m\in{\mathbb{Z}},\ k\in{\mathbb{N}}\right\}.$
There, we have some additional bounds:
Theorem 6.6.
If
$\beta\gamma_{A}<\gamma_{B}$
and
$\frac{1-\alpha\beta}{\alpha\gamma_{B}}\geq\frac{9\beta}{1-\beta}$
and
$\alpha\beta>\frac{1}{7},$
then the set $\widetilde{\text{3-BA}}(1/6)$ is $(\alpha,\beta)$-losing.
Proof.
Let
$2\varepsilon={\gamma_{B}}-\beta\gamma_{A}.$
Then, Bob will play in the following way: on his first turn, he plays an interval of length
$|B_{0}|=\frac{1-\alpha\beta}{\alpha\gamma_{B}}.$
Then, he will play arbitrarily until some turn $t$, where
$\frac{1-\alpha\beta}{\alpha\gamma_{B}}\cdot\frac{1}{3^{n+2}}\leq|B_{t}|<(1+\varepsilon)\cdot\frac{1-\alpha\beta}{\alpha\gamma_{B}}\cdot\frac{1}{3^{n+2}}<\frac{1-\alpha\beta}{\alpha\beta\gamma_{A}}\cdot\frac{1}{3^{n+2}}$
for some $n\in\mathbb{N}$. Then, on turn $t$, he will center his ball at a $c_{t}$ such that
$c_{t}=\frac{m}{3^{n}}$
for some $m\in\mathbb{Z}$. On his next move, Bob is always able to play in one of the following two ways:
  • Bob will center his next interval such that
    $\left|c_{t+1}-\frac{m}{3^{n}}\right|<\frac{1}{6\cdot 3^{n}}-\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}$ (6.3)
    and Bob will try to minimize $|c_{t+1}-\frac{m}{3^{n}}|$, or
  • Bob will center his next interval such that
    $\left|c_{t+1}-\left(\frac{m}{3^{n}}\pm\frac{1}{3^{n+1}}\right)\right|<\frac{1}{6\cdot 3^{n+1}}-\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}$ (6.4)
    and Bob will try to minimize $|c_{t+1}-\frac{m}{3^{n}}\mp\frac{1}{3^{n}+1}|$.
We show that this is always possible. First, we show that Bob can play his strategy on turn $t$. The length condition is equivalent to
$0\leq\left\{\log_{3}\frac{|B_{0}|\alpha\gamma_{B}}{1-\alpha\beta}+t\log_{3}\alpha\beta\right\}<\log_{3}(1+\varepsilon)$
and by the Equidistribution Theorem, there are infinitely many values of $t$ that satisfy the inequality. Then, we have
$|B_{t}|\geq\frac{1-\alpha\beta}{\alpha\beta\gamma_{A}}\cdot\frac{1}{3^{n+2}}\geq\frac{\beta}{1-\beta}\cdot\frac{1}{3^{n}}$
so $(1-\beta)|A_{t-1}|\geq 1/3^{n}$ and therefore Bob can always play an interval of the form $m/3^{n}$ on turn $t$.
Now, we show Bob can play his strategy on turn $t+1$. First, we note that since $\alpha\beta>1/7$, we have
$\frac{1}{6\alpha\beta}$ $<\frac{7}{6}$
$\frac{|B_{t}|\gamma_{A}}{2(1-\alpha\beta)}<\frac{1}{2\alpha\beta}\cdot\frac{1}{3^{n+2}}$ $<\frac{7}{6\cdot 3^{n+1}}$
$\frac{|B_{t}|\gamma_{A}}{2}+\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}$ $<\frac{1}{3^{n+1}}+\frac{1}{6\cdot 3^{n+1}}$
$\frac{m}{3^{n}}+\frac{|B_{t}|\gamma_{A}}{2}$ $<\frac{m}{3^{n}}+\frac{1}{3^{n+1}}+\frac{1}{6\cdot 3^{n+1}}-\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)},$
which implies that Bob is able to ensure that $c_{t+1}<\frac{m}{3^{n}}+\frac{1}{3^{n+1}}+\frac{1}{6\cdot 3^{n+1}}-\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}$. By similar logic, Bob can ensure that $c_{t+1}>\frac{m}{3^{n}}-\frac{1}{3^{n+1}}-\frac{1}{6\cdot 3^{n+1}}+\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}$. Furthermore, we note that
$\frac{|B_{t}|\alpha\gamma_{B}}{1-\alpha\beta}$ $>\frac{1}{3^{n+2}}$
$|B_{t}|\alpha\left(1-\beta-\frac{\beta\gamma_{A}}{1-\alpha\beta}\right)$ $>\frac{1}{3^{n+2}}$
$\frac{1}{3^{n+2}}+\frac{|B_{t}|\alpha\beta\gamma_{A}}{1-\alpha\beta}$ $<(1-\beta)\alpha|B_{t}|$
$\left|\left(\frac{m}{3^{n}}\pm\left(\frac{1-\frac{1}{6}}{3^{n+1}}+\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}\right)\right)-\left(\frac{m}{3^{n}}\pm\left(\frac{1}{6\cdot 3^{n}}-\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}\right)\right)\right|$ $<(1-\beta)\alpha|B_{t}|,$
which implies that Bob is always able to satisfy either 6.3 or 6.4.
Finally, we show this strategy is winning. This follows from Lemma 6.3: if 6.3 is true, then in finitely many $k$ turns, $|B_{t+k}|\subset(\frac{m-1/6}{3^{n}},\frac{m+1/6}{3^{n}})$. Since Bob can play this way for infinitely many sets of two consecutive turns, Bob wins. ∎
Conjecture 6.7.
If
$\beta\gamma_{A}\geq\gamma_{B},$
then the set $\widetilde{\text{3-BA}}(1/6)$ is $(\alpha,\beta)$-winning. If not, then it is $(\alpha,\beta)$-losing.
@@ Acknowledgements
We are very grateful to Prof. Dmitry Kleinbock for the project proposal and his valuable suggestions in research paths. We would also like to thank Prof. Pavel Etingof, Dr. Slava Gerovitch, Dr. Felix Gotti, Dr. Tanya Khovanova, and all the PRIMES organizers for providing the opportunity to conduct math research through this program.

### v3

@@ 5. Proofs of the statements regarding d-Badly Approximable Numbers
    $\left(\frac{1}{1-\beta}\right)\leq d^{N}|A_{t}|\leq\left(\frac{1}{1-\beta}\right)(1+\varepsilon)=\left(\frac{2c(1-\alpha\beta)}{\beta\gamma}\right),$ (5.2)
    and since $|A_{t}|(1-\beta)>1/d^{N}$, Bob can play his next interval centered at $m/d^{N}$ for some $m\in\mathbb{Z}$. Furthermore, by Lemma 1.2, Bob can play in such a way that in finitely many moves, $B_{t+s}\subset J_{n}$. By Lemma 4.1, there exists an infinite number of possible $t$ that satisfy inequality 5.2. Therefore, Bob can do this an infinite number of times, and Alice loses.
Proof of Theorem 2.6 We will describe the winning strategy for Alice. Suppose on the step $t$ of the game Bob played an interval $B_{t}$ of length $l_{t}$. Pick $k_{t}$ such that
$\frac{l_{t}}{2}\leq\frac{1}{2^{k_{t}}}<l_{t}\leq\frac{1}{2^{k_{t}-1}}.$ (5.3)
Alice’s strategy is to choose a ball of length $\alpha l_{t}$ and center $c_{t}$ of the form $(r\pm\frac{1}{3})2^{-k_{t}}$ for some $r\in{\mathbb{Z}}$ as $A_{t}$. She can always do this, as for any two subsequent points $c_{t}^{1},c_{t}^{2}$ of the chosen form one has
$\left|c^{1}_{t}-c^{2}_{t}\right|+\alpha l_{t}=\frac{1}{3\cdot 2^{k_{t}-1}}+\alpha l_{t}<\frac{2l_{t}}{3}+\frac{l_{t}}{3}=l_{t}.$
In some special cases, when needed, we will specify which of such balls Alice will need to pick.
Define $q_{t}$ to be the smallest integer such that $m/2^{q_{t}}\in A_{t}$ for some $m\in{\mathbb{Z}}$. Note that $q_{t}>k_{t}$, as no points of the form $\frac{s}{2^{k_{t}}}$ with $s\in\mathbb{Z}$ belong to $A_{t}$ by construction. One has
$\left|\frac{m}{2^{q_{t}}}-c_{t}\right|=\frac{1}{3\cdot 2^{q_{t}}}<\frac{\alpha l_{t}}{2}<\frac{1}{3\cdot 2^{q_{t}-1}}.$ (5.4)
@@
$\left|x-\frac{m}{2^{n}}\right|>\frac{c}{2^{n}}\ \text{for all}\ n\leq q_{i}-2$
by induction on the number of step $i$. Suppose that the statement holds for $i$, and Bob picks an interval $B_{i+1}$ with length $l_{i+1}=\alpha\beta l_{i}$.
  • First, we consider $q_{i}\leq n\leq k_{i+1}$. Note that for all $x\in A_{i+1}$ and any $m\in\mathbb{Z}$
@@
    $>\frac{c}{2^{q_{i}}}\geq\frac{c}{2^{n}},$
    where we used 5.4 in the second to last inequality.
  • To take care of $k_{i+1}+1\leq n\leq q_{i+1}-2$, simply note that for any $x\in A_{i+1}$ and any $c<1/6$,
    $\left|x-\frac{m}{2^{n}}\right|$ $>\frac{1}{3\cdot 2^{n}}-\frac{\alpha l_{i+1}}{2}>\frac{1}{3\cdot 2^{n}}-\frac{1}{3\cdot 2^{q_{i+1}-1}}>\frac{1}{6\cdot 2^{n}}\geq\frac{c}{2^{n}}.$
    Here, we again used 5.4 in the second inequality.
  • Finally, consider $n=q_{i}-1$. There are two possible cases based on $A_{i}$, the prior move.
@@
      then for all $x\in A_{i}$ and $m\in\mathbb{Z}$
      $\left|x-\frac{m}{2^{q_{i}-1}}\right|$ $\geq\frac{1}{3\cdot 2^{q_{i}-1}}-\frac{\alpha l_{i}}{2}>\frac{c}{2^{q_{i}-1}},$
      which means that the desired inequality is already satisfied for any $x\in A_{i}$ and thus for any $x\in A_{i+1}$.
@@
      If
      $B_{i+1}\cap\left[\frac{m-c}{2^{q_{i}-1}},\frac{m+c}{2^{q_{i}-1}}\right]=\emptyset$
      for any $m\in\mathbb{Z}$, it is clear that Alice does not have to worry about $n=q_{i}-1$, since for all $x\in A_{i+1}\subset B_{i+1}$, $|x-m/2^{q_{i}-1}|>c/2^{q_{i}+1}$. Thus, for the remainder of this case, assume $B_{i+1}$ intersects such an interval. Without loss of generality (the other case is symmetric), assume also that $m/2^{q_{i}-1}>x$ for all $x\in B_{i+1}$ (that is, the fraction lies to the right of the interval $B_{i+1}$). Alice will then choose an interval $A_{i+1}$ according to her strategy which is centered at $c_{i+1}$ such that
@@
      that is, Alice can not choose the closest interval to $\frac{m}{2^{q_{i}-1}}$. This is always possible because the second closest interval of the desired form is within $l_{i+1}$ of $\frac{m}{2^{q_{i}-1}}$. Indeed: the left endpoint of the second closest interval of this form is
      $e_{i+1}=\frac{m}{2^{q_{i}-1}}-\frac{2}{3}\left(\frac{1}{2^{k_{i+1}}}\right)-\frac{\alpha l_{i+1}}{2},$
      while the left endpoint of $B_{i+1}$ is smaller than $\frac{m}{2^{q_{i}-1}}-l_{i+1}<e_{i}$, where the latter inequality holds as
      $\frac{2}{3}\left(\frac{1}{2^{k_{i+1}}}\right)+\frac{\alpha l_{i+1}}{2}<\frac{2}{3}l_{i+1}+\frac{1}{6}l_{i+1}<l_{i+1}$
      by 5.3.
      In this case, for $x\in A_{i+1}$ one has
@@
      $>\frac{c}{2^{q_{i}-1}},$
      proving the desired inequality in this last case. This completes the induction, and thus the whole proof.
@@ Acknowledgements
We are very grateful to Prof. Dmitry Kleinbock for the project proposal and his valuable suggestions in research paths. We would like to thank Prof. Lior Fishman and Prof. Stephen Jackson for fruitful discussions of the Schmidt game determinacy problems. We would also like to thank Prof. Pavel Etingof, Dr. Slava Gerovitch, Dr. Felix Gotti, Dr. Tanya Khovanova, and all the PRIMES organizers for providing the opportunity to conduct math research through this program.